OpenStudy (anonymous):

i neeed help!!!!! Solve the trigonometric equation.. sinx tan^2x=sinx

OpenStudy (perl):

your real name is shelly poop?

OpenStudy (anonymous):

no. its shelly.

OpenStudy (shubhamsrg):

then why poop? o.O

OpenStudy (anonymous):

i just randomly made up a username

OpenStudy (shubhamsrg):

quite some name that is! o.O

OpenStudy (shubhamsrg):

anyways,,for the question,,bring everything in LHS..what do you get?

OpenStudy (anonymous):

thanks . haha LHS?? whats that?

OpenStudy (anonymous):

left hand side :P

OpenStudy (anonymous):

i know i have to equal it to zero , i think , or put it all on one side?

OpenStudy (anonymous):

ohhh haha i knew that >.<

OpenStudy (shubhamsrg):

yep./..put all on one side..

OpenStudy (shubhamsrg):

now take sinx common out..what do you get ?

OpenStudy (anonymous):

\[sinx \tan^2x=sinx .\] That's the question, no?

OpenStudy (anonymous):

yeahh thats the question .. i dont make tan^2 to 1-sec^2x ??

hartnn (hartnn):

i think that is not required, if u know when tan x =0 first, bring all terms on one side, what u get ?

OpenStudy (anonymous):

you get sinx tan^2x -sinx .. then the sin cancels out? nd tan^2x=0 ?

hartnn (hartnn):

its not accurate to say that sin cancels out you factor out sin x like this : sinx (tan^2x-1) = 0 then u equate both factors to 0.

OpenStudy (anonymous):

oohh okay thank you very much!:]

hartnn (hartnn):

so , can u solve for x in sinx (tan^2x-1) = 0 ?

OpenStudy (anonymous):

umm i think.. haha ill try to figure it out >.<

hartnn (hartnn):

sure, if u don't get it, ask....

OpenStudy (anonymous):

do i move the 2 in front of the tan? haha. >.<

hartnn (hartnn):

sinx (tan^2x-1) = 0 equating both factors to 0, u get sin x=0 tan^2 x = 1 right ?

OpenStudy (anonymous):

yeaaahh

hartnn (hartnn):

so what is x, when sin x=0 ??

OpenStudy (anonymous):

pi and 2pi ,, sorry i dont have the little symbol

hartnn (hartnn):

yeah, do u have any interval to solve in ?? or u want general solution ??

OpenStudy (anonymous):

general

hartnn (hartnn):

so, starting with 0, sin x is 0 for x=0,\(\pi\),2\(\pi\),3\(\pi\),........n\(\pi\) got this ?

OpenStudy (anonymous):

yeah i understand that part ,

OpenStudy (anonymous):

the tan part i dont get

hartnn (hartnn):

\(\tan^2x = 1\) so, tan x = \(\pm\)1 right ? u know for which values of x are tan values + or -1

OpenStudy (anonymous):

what happened to the ^2 ?

hartnn (hartnn):

taking square root on both sides....

OpenStudy (anonymous):

ohhhhhhh you squared it! . ahh i feel so dumb-.- , ahh yess thank you so much!! i know the values now :] .

hartnn (hartnn):

great :) welcome ^_^

OpenStudy (anonymous):

thanks again!!:D